# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ B.1.3 Acceptance probability
$=\exp\bigg{(}-{1\over 2\sigma^{2}}\sum_{i=1}^{n}\bigg{(}\mathfrak{a}_{\mathbf{v}_{k,i}}\beta_{k}(z_{k}-z_{k}^{\text{new}})\bigg{(}2y_{i}-\sum_{t=1}^{T_{\max}}z_{t}\mathfrak{a}_{\mathbf{v}_{t,i}}\beta_{t}-\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}\mathfrak{a}_{\mathbf{v}_{t,i}}\beta_{t}\bigg{)}\bigg{)}\bigg{)}\times{{T_{\max}\choose\sum_{t=1}^{T_{\max}}z_{t}}\over{T_{\max}\choose\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}}}$.
Finally, we accept $\mathbf{z}^{\text{new}}$ with a acceptance probability $\tau_{\text{accept}}$, where
$\tau_{\text{accept}}=\min\left(1,{\pi\{\mathbf{z}^{\text{new}}|y_{1},...,y_{n},\text{others}\}q(\mathbf{z}|\mathbf{z}^{\text{new}})\over\pi\{\mathbf{z}|y_{1},...,y_{n},\text{others}\}q(\mathbf{z}^{\text{new}}|\mathbf{z})}\right).$
Appendix C Proofs of Theorem 5.1
C.1 Additional notations
@@ C.2 Overall strategy
Here, we let $A^{(1)}=A$ and $A^{(-1)}=A^{c}$ for a given set $A$. For a given binary-product partition $\mathcal{R},$ let $\text{var}(\mathcal{R})$ and $\text{sval}(\mathcal{R})$ be the set of split variables and the set of split values used in constructing $\mathcal{R},$ respectively. Since there is an one-to-one relation between $(S,\mathbf{s})$ and $\mathcal{R},$ we parameterize a given identifiable binary-product tree by $\mathcal{R},\beta$ instead of $S,\mathbf{s},\beta.$ We will use these two parameterizations interchangeably unless there is any confusion.
@@
$\lim_{n\to\infty}\mathbb{P}_{0}^{n}\bigg{\{}\pi_{\xi}\{\|f-f_{0}\|_{2,n}+|\sigma^{2}-\sigma_{0}^{2}|>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\bigg{\}}=0.$ (C.1)
We prove (C.1) as follows. We first specify a subset $A_{n}$ of $\mathcal{X}^{n}$ such that $\mathbb{P}_{\mathbf{X}}^{n}\{A_{n}\}\rightarrow 1$ as $n\rightarrow\infty,$ and for any $\mathbf{x}^{(n)}\in A_{n}$ we have
$\mathbb{P}_{Y^{(n)}}\bigg{\{}\{\pi_{\xi}\{\|f-f_{0}\|_{2,n}+|\sigma^{2}-\sigma_{0}^{2}|>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\}\bigg{|}\mathbf{X}^{(n)}=\mathbf{x}^{(n)}\bigg{\}}\rightarrow 0$ (C.2)
@@
$\mathbb{P}_{0}^{n}\bigg{\{}\{\pi_{\xi}\{\|f-f_{0}\|_{2,n}+|\sigma^{2}-\sigma_{0}^{2}|>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\}\bigg{\}}$
$\leq\mathbb{P}_{0}^{n}\bigg{\{}\{\pi_{\xi}\{\|f-f_{0}\|_{2,n}+|\sigma^{2}-\sigma_{0}^{2}|>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\}\cap A_{n}\bigg{\}}+\mathbb{P}_{\mathbf{X}}^{n}(A_{n}^{c})$
@@
Let $C_{\mathbb{B}}$ and $d_{\sigma}$ be positive constants to be determined later (Section C.7 of Supplementary Material). To prove (C.2), we verify the following three conditions for $\mathbf{x}^{(n)}\in A_{n}:$ there exists $\mathcal{F}^{n}\subset\mathcal{F}$ depending on $\mathbf{x}^{(n)}$ such that
$\log N\left({\epsilon_{n}\over 72},\mathcal{F}^{n},\|\cdot\|_{\infty}\right)\leq n\epsilon_{n}^{2}$ (C.3)
$\pi\big{\{}f\in\mathcal{F}:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}\big{\}}\geq e^{-d_{1}n\epsilon_{n}^{2}}$ (C.4)
$\pi\{\mathcal{F}\backslash\mathcal{F}^{n}\}\leq e^{-(2d_{1}+d_{\sigma}+2)n\epsilon_{n}^{2}}$ (C.5)
for some constant $d_{1}>0$. In turn, we show that these three conditions imply the posterior convergence rate in (C.2) by checking the conditions in Theorem 4 of nonp1 .
C.3 Construction of $A_{n}$
For constructing $A_{n},$ we need the following theorem which shows that any populationally identifiable smooth function can be approximated by a sum of identifiable binary-product trees whose heights are bounded. This theorem is technically involved, and its proof is given in Section F of Supplementary Material.
Theorem C.1.
For $\alpha\in(0,1]$, suppose $g_{S}$ be a populationally identifiable $S$-component function in $\mathcal{H}^{\alpha}_{F}$. Then, there exist positive constants $m$ and $C$ such that for any $T_{S}\in\mathbb{N}_{+},$ there exist $T_{S}$ many binary-product partitions $\mathcal{R}_{1},\ldots,\mathcal{R}_{T_{S}}$ with $var(\mathcal{R}_{t})=S$ for all $t\in[T_{S}]$ and $T_{S}$ many real numbers $\beta_{1},\ldots,\beta_{T_{S}}$ depending on $\mathbf{x}^{(n)}$ such that $\sup_{j}|\beta_{j}|\leq C$ and
$\left\|g_{S}(\cdot)-\sum_{t=1}^{T_{S}}\mathbb{T}(\cdot:\mathcal{R}_{t},\beta_{t})\right\|_{\infty}\leq\|g_{S}\|_{\mathcal{H}^{\alpha}}{m^{\alpha}\over(T_{S}^{1\over|S|}+1)^{\alpha}}+\Phi_{n,S}(\mathbf{x}^{(n)},T_{S}),$ (C.6)
@@ C.3 Construction of $A_{n}$
for all sufficiently large $n.$
Let $t_{n,S}$ be the smallest positive integer satisfying
$\|f_{0,S}\|_{\mathcal{H}^{\alpha}}{m^{\alpha}\over(t_{n,S}^{1\over|S|}+1)^{\alpha}}+(24F+1)2^{|S|}\sqrt{(t_{n,S}^{1\over|S|}+1)^{|S|}}\sqrt{\log n\over n}\leq{C_{\mathbb{B}}\epsilon_{n}\over 2^{p+1}}.$ (C.7)
@@
$A_{n}:=\left\{\mathbf{x}^{(n)}:\Phi_{n,S}(\mathbf{x}^{(n)},t_{n,S})\leq(24F+1)2^{|S|}\sqrt{(t_{n,S}^{1\over|S|}+1)^{|S|}}\sqrt{\log n\over n}\mbox{ for all }S\subseteq[p]\right\}.$ (C.8)
By Theorem C.1, $\mathbb{P}_{\mathbf{X}}^{n}(A_{n})\rightarrow 1$ as $n\rightarrow\infty.$
C.4 Proof of Condition (C.3)
@@
for some positive integer $M_{n}\leq T_{\max},$ and will choose $M_{n}$ so that $\mathcal{F}^{n}$ satisfies Condition (C.3) with $\epsilon_{n}^{2}=n^{-{2\alpha\over 2\alpha+p}}\log n$. For this purpose, we need the following lemma.
Lemma C.2.
@@
$N(\epsilon_{1},[-n,n]^{T},\|\cdot\|_{1})\leq\bigg{(}1+{2nT\over\epsilon_{1}}\bigg{)}^{T}$ (C.9)
@@
$\leq\sum_{t=1}^{T}\sup_{\mathbb{v}\in\{-1,1\}^{|S_{t}|}}|\beta_{t,\mathbb{v}}-\tilde{\beta}_{t,\mathbb{v}}|$
$=\sum_{t=1}^{T}\sup_{\mathbb{v}\in\{-1,1\}^{|S_{t}|}}|\mathfrak{a}_{t,\mathbf{v}}||\beta_{t}-\tilde{\beta}_{t}|$
$\leq\sum_{t=1}^{T}n^{|S_{t}|}|\beta_{t}-\tilde{\beta}_{t}|$
$\leq n^{p}\epsilon_{1},$
where we use the inequality that $\sup_{\mathbb{v}\in\{-1,1\}^{|S_{t}|}}|\mathfrak{a}_{t,\mathbf{v}}|\leq n^{|S_{t}|}$ obtained in Proposition A.2. By letting $\epsilon_{1}=\epsilon/n^{p}$ in (C.9), we have
$N(\epsilon,\mathcal{F}^{n}(T,\mathcal{E}),\|\cdot\|_{\infty})\leq\bigg{(}1+{2n^{p+1}T\over\epsilon}\bigg{)}^{T},$ (C.10)
@@
$N\bigg{(}{\epsilon_{n}\over 72},\mathcal{F}^{n},\|\cdot\|_{\infty}\bigg{)}$ $\leq\sum_{T=1}^{M_{n}}\sum_{\mathcal{E}\in\mathcal{E}(T)}N\bigg{(}{\epsilon_{n}\over 72},\mathcal{F}^{n}(T,\mathcal{E}),\|\cdot\|_{\infty}\bigg{)}$
$\leq\sum_{T=1}^{M_{n}}\bigg{(}1+{72n^{p+1}T\over\epsilon_{n}}\bigg{)}^{T}\sum_{\mathcal{E}\in\mathcal{E}(T)}1$
$\lesssim\sum_{T=1}^{M_{n}}\bigg{(}1+{72n^{p+1}T\over\epsilon_{n}}\bigg{)}^{T}n^{Tp}$ (C.11)
$\leq M_{n}n^{M_{n}p}\bigg{(}1+{72n^{p+1}M_{n}\over\epsilon_{n}}\bigg{)}^{M_{n}},$
where the inequality in (C.11) is due to
@@
$\log N\bigg{(}\epsilon_{n}^{2},\mathcal{F}^{n},\|\cdot\|_{\infty}\bigg{)}\lesssim\log M_{n}+M_{n}p\log n+M_{n}\log\bigg{(}1+{72n^{p+1}M_{n}\over\epsilon_{n}}\bigg{)}$
With $M_{n}=\lfloor C_{1}{n\epsilon_{n}^{2}\over\log n}\rfloor$ for some large enough constant $C_{1}>0$ (will be determined later), Condition (C.3) is satisfied.
C.5 Proof of Condition (C.4)
For given $\mathbf{x}^{(n)}\in A_{n},$ let $\hat{f}_{S}$ be an ensemble of $t_{n,S}$ many identifiable binary-product trees satisfying $\|f_{0,S}-\hat{f}_{S}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}/2^{p+1},$ whose existence is guaranteed by Theorem C.1, and let $\hat{f}=\sum_{S\subseteq[p]}\hat{f}_{S}.$ Let $\hat{\mathcal{E}}_{S}$ and $\hat{\mathcal{B}}_{S}$ be the ensemble partitions and height vector for $\hat{f}_{S}$ and let $\hat{\mathcal{E}}$ and $\hat{\mathcal{B}}$ be the ensemble partition and height vector for $\hat{f}$. Let $t_{n}=\sum_{S\subseteq[p]}t_{n,S}$.
@@
$\leq\sum_{t=1}^{t_{n}}n^{p}|\beta_{t}-\hat{\beta}_{t}|$
$\leq n^{p}\sqrt{t_{n}}\|\mathcal{B}-\hat{\mathcal{B}}\|_{2},$
where the second inequality follows from Proposition A.2. Hence, if $\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq(n^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2,$ we have
@@
Thus, to prove Condition (C.4), it suffices to show that
$\pi\left\{T=t_{n},\mathcal{E}=\hat{\mathcal{E}},\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq(n^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2\right\}$ (C.12)
is sufficiently large. We decompose (C.12) as
$\pi\left\{T=t_{n},\mathcal{E}=\hat{\mathcal{E}},\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq(n^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2\right\}$
$=\pi\{T=t_{n}\}\pi\{\mathcal{E}=\hat{\mathcal{E}}|T=t_{n}\}\pi\{\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq(n^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2|\mathcal{E}=\hat{\mathcal{E}},T=t_{n}\}.$
@@
C-(a). A lower bound of $\pi\{T=t_{n}\}:$ Since $t_{n}\lesssim n\epsilon_{n}^{2}/\log n$, there exists a constant $d_{2}>0$ such that
$\pi\{T=t_{n}\}$ $={\exp(-C_{*}t_{n}\log n)\over\sum_{t=0}^{T_{\max}}\exp(-C_{*}t_{n}\log n)}$
@@
C-(b). A lower bound of $\pi\{\mathcal{E}=\hat{\mathcal{E}}|T=t_{n}\}:$ Let $\mathcal{E}=(\mathcal{R}_{t},t=1,\ldots,t_{n})$ and $\hat{\mathcal{E}}=(\hat{\mathcal{R}}_{t},t=1,\ldots,t_{n}).$ Note that conditional on $T=t_{n},$ $\mathcal{R}_{t},t\in[t_{n}]$ are independent a priori and thus we have
@@
for a certain constant $d_{3}>0.$ C-(c). A lower bound of $\pi\{\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq(n^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2|\mathcal{E}=\hat{\mathcal{E}},T=t_{n}\}:$
$\pi\left\{\mathcal{B}\in\mathbb{R}^{t_{n}}:\|\hat{\mathcal{B}}-\mathcal{B}\|_{2}\leq{C_{\mathbb{B}}\over n^{p}\sqrt{t_{n}}}\times{\epsilon_{n}\over 2}\right\}$
$\geq 2^{-t_{n}}\left({t_{n}\over 2}\right)^{-{t_{n}\over 2}-1}\exp\left(-{\|\hat{\mathcal{B}}\|_{2}^{2}\over\sigma_{\beta}^{2}}-{(C_{\mathbb{B}}\epsilon_{n})^{2}\over 8n^{2p}t_{n}\sigma^{2}_{\beta}}\right)\left({(C_{\mathbb{B}}\epsilon_{n})^{2}\over 4n^{2p}t_{n}\sigma_{\beta}^{2}}\right)^{t_{n}\over 2}$ (C.13)
$\geq 2^{-t_{n}}\left({t_{n}\over 2}\right)^{-{t_{n}\over 2}-1}\exp\left(-{t_{n}C^{2}\over\sigma_{\beta}^{2}}-{(C_{\mathbb{B}}\epsilon_{n})^{2}\over 8n^{2p}t_{n}\sigma_{\beta}^{2}}\right)\left({(C_{\mathbb{B}}\epsilon_{n})^{2}\over 4n^{2p}t_{n}\sigma_{\beta}^{2}}\right)^{t_{n}\over 2},$ (C.14)
where (C.13) is derived from equation (8.9) in (bforest ) and (C.14) is derived from Theorem C.1. Finally, the three terms in (C.14) are bounded below by
@@
$\exp\left(-{t_{n}C^{2}\over\sigma_{\beta}^{2}}-{(C_{\mathbb{B}}\epsilon_{n})^{2}\over 8n^{2p}t_{n}\sigma_{\beta}^{2}}\right)$ $\geq\exp(-d_{5}n\epsilon_{n}^{2})$ (C.15)

### v3

@@ A.3.1 Proposal for $\beta_{t}$
${\partial\over\partial\beta_{t}}\log\pi\{\beta_{t}|\text{others}\}$ $={\partial\over\partial\beta_{t}}\mathcal{L}(T,\{S_{k},\mathbf{s}_{k},\beta_{k}\}_{k=1}^{T},\psi)-{\beta_{t}\over\sigma_{\beta}^{2}}.$
A.3.2 Transition probability ratio
The transition probability ratio for the proposal distribution is given as
${q(\beta_{t}|\beta_{t}^{\text{new}})\over q(\beta_{t}^{\text{new}}|\beta_{t})}=\exp\bigg(-{1\over 2}\big((m^{\text{new}})^{2}-m^{2}\big)\bigg),$
where
$m^{\text{new}}=m+{\varepsilon\over 2}{\partial\over\partial\beta_{t}}\log\pi\{\beta_{t}|\text{others}\}+{\varepsilon\over 2}{\partial\over\partial\beta_{t}}\log\pi\{\beta_{t}^{\text{new}}|\text{others}\}$
and $\varepsilon$ is the step size.
A.3.3 Posterior probability ratio
Since
${\pi\{\beta_{t}^{\text{new}}\}\over\pi\{\beta_{t}\}}=\exp\bigg(-{1\over 2\sigma_{\beta}^{2}}(\beta_{t}^{\text{new}}-\beta_{t})^{2}\bigg),$
the posterior probability ratio is
${\pi\{\beta_{t}^{\text{new}}|\text{others}\}\over\pi\{\beta_{t}|\text{others}\}}={\mathcal{L}(T,\{S_{k},\mathbf{s}_{k},\beta_{k}\}_{k\neq t},S_{t},\beta_{t}^{\text{new}},\mathbf{s}_{t}^{\text{new}},\psi)\over\mathcal{L}(T,\{S_{k},\mathbf{s}_{k},\beta_{k}\}_{k=1}^{T},\psi)}\exp\bigg(-{1\over 2\sigma_{\beta}^{2}}(\beta_{t}^{\text{new}}-\beta_{t})^{2}\bigg).$
A.3.4 Acceptance probability
In summary, we accept $\beta_{t}^{\text{new}}$ with probability $p_{\text{accept}},$ where
$p_{\text{accept}}=\min\left\{1,{\pi\{\beta_{t}^{\text{new}}|\text{others}\}\over\pi\{\beta_{t}|\text{others}\}}{q(\beta_{t}|\beta_{t}^{\rm new})\over q(\beta_{t}^{\rm new}|\beta_{t})}\right\}.$
A.4 Sampling nuisance parameter $\psi$
In this section, we consider the gaussian regression model, i.e., $y_{i}$ is realization from
$Y_{i}|\mathbf{x}_{i}\sim N(\cdot|f(\mathbf{x}_{i}),\sigma^{2})$
for i=1,,ni=1,...,n, where $\sigma^{2}$ is the nuisance parameter. Since $\sigma^{2}\sim IG\left({v\over 2},{v\lambda\over 2}\right)$, it follows that
$\sigma^{2}|\text{others}\sim IG\bigg({v+n\over 2},{v\lambda+\sum_{i=1}^{n}(y_{i}-f(\mathbf{x}_{i}))^{2}\over 2}\bigg).$
Appendix B Proofs of Theorem 5.1
B.1 Additional notations
@@ B.2 Overall strategy
Here, we let $A^{(1)}=A$ and $A^{(-1)}=A^{c}$ for a given set $A$. For a given binary-product partition $\mathcal{R},$ let $\text{var}(\mathcal{R})$ and $\text{sval}(\mathcal{R})$ be the set of split variables and the set of split values used in constructing $\mathcal{R},$ respectively. Since there is a one-to-one relation between $(S,\mathbf{s})$ and $\mathcal{R},$ we parameterize a given identifiable binary-product tree by $\mathcal{R},\beta$ instead of $S,\mathbf{s},\beta.$ We will use these two parameterizations interchangeably unless there is any confusion.
@@
$\lim_{n\to\infty}\mathbb{P}_{0}^{n}\bigg\{\pi_{\xi}\{\|f-f_{0}\|_{2,n}>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\bigg\}=0.$ (B.1)
We prove (B.1) as follows. We first specify a subset $A_{n}$ of $\mathcal{X}^{n}$ such that $\mathbb{P}_{\mathbf{X}}^{n}\{A_{n}\}\rightarrow 1$ as $n\rightarrow\infty,$ and for any $\mathbf{x}^{(n)}\in A_{n}$ we have
$\mathbb{P}_{Y^{(n)}}\bigg\{\{\pi_{\xi}\{\|f-f_{0}\|_{2,n}>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\}\bigg|\mathbf{X}^{(n)}=\mathbf{x}^{(n)}\bigg\}\rightarrow 0$ (B.2)
@@
$\mathbb{P}_{0}^{n}\bigg\{\{\pi_{\xi}\{\|f-f_{0}\|_{2,n}>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\}\bigg\}$
$\leq\mathbb{P}_{0}^{n}\bigg\{\{\pi_{\xi}\{\|f-f_{0}\|_{2,n}>B_{n}\epsilon_{n}|\mathbf{X}^{(n)},Y^{(n)}\}>\delta\}\cap A_{n}\bigg\}+\mathbb{P}_{\mathbf{X}}^{n}(A_{n}^{c})$
@@
Let $C_{\mathbb{B}}$ be a positive constant to be determined later (Section B.7 of Supplementary Material). To prove (B.2), we verify the following three conditions for $\mathbf{x}^{(n)}\in A_{n}:$ there exists $\mathcal{F}^{n}\subseteq\mathcal{F}$ depending on $\mathbf{x}^{(n)}$ such that
$\log N\left({\epsilon_{n}\over 36},\mathcal{F}^{n},\|\cdot\|_{\infty}\right)\leq n\epsilon_{n}^{2}$ (B.3)
$\pi\big\{f\in\mathcal{F}:\|f-f_{0}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}\big\}\geq e^{-d_{1}n\epsilon_{n}^{2}}$ (B.4)
$\pi\{\mathcal{F}\backslash\mathcal{F}^{n}\}\leq e^{-(2d_{1}+2)n\epsilon_{n}^{2}}$ (B.5)
for some constant $d_{1}>0$. In turn, we show that these three conditions imply the posterior convergence rate in (B.2) using Theorem 4 of Ghosal and van der Vaart (2007).
B.3 Construction of $A_{n}$
For constructing $A_{n},$ we need the following theorem which shows that any populationally identifiable smooth function can be approximated by a sum of identifiable binary-product trees whose heights are bounded. This theorem is technically involved, and its proof is given in Section E of Supplementary Material.
Theorem B.1.
For $\alpha\in(0,1]$, suppose $g_{S}$ be a populationally identifiable $S$-component function in $\mathcal{H}^{\alpha}_{F}$. Then, there exist positive constant $C$ that does not depend on the function $g_{S}$ such that for any $T_{S}\in\mathbb{N}_{+},$ there exist $T_{S}$ many binary-product partitions $\mathcal{R}_{1},\ldots,\mathcal{R}_{T_{S}}$ with $var(\mathcal{R}_{t})=S$ for all $t\in[T_{S}]$ and $T_{S}$ many real numbers $\beta_{1},\ldots,\beta_{T_{S}}$ depending on $\mathbf{x}^{(n)}$ such that $\sup_{j}|\beta_{j}|\leq C$ and
$\left\|g_{S}(\cdot)-\sum_{t=1}^{T_{S}}\mathbb{T}(\cdot:\mathcal{R}_{t},\beta_{t})\right\|_{\infty}\leq\|g_{S}\|_{\mathcal{H}^{\alpha}}{C_{S}^{\alpha}\over(T_{S}^{1\over|S|}+1)^{\alpha}}+\Phi_{n,S}(\mathbf{x}^{(n)},T_{S}),$ (B.6)
@@ B.3 Construction of $A_{n}$
for all sufficiently large $n$ and $C_{S}=\sqrt{|S|}/p_{L}$.
For $S\in\mathbb{S}$, let $t_{n,S}$ be the smallest positive integer satisfying
$\|f_{0,S}\|_{\mathcal{H}^{\alpha}}{C_{S}^{\alpha}\over(t_{n,S}^{1\over|S|}+1)^{\alpha}}+(24F+1)2^{|S|}\sqrt{(t_{n,S}^{1\over|S|}+1)^{|S|}}\sqrt{\log n\over n}\leq{C_{\mathbb{B}}\epsilon_{n}\over 2^{p+1}}.$ (B.7)
@@
$A_{n}:=\left\{\mathbf{x}^{(n)}:\Phi_{n,S}(\mathbf{x}^{(n)},t_{n,S})\leq(24F+1)2^{|S|}\sqrt{(t_{n,S}^{1\over|S|}+1)^{|S|}}\sqrt{\log n\over n}\mbox{ for all }S\in\mathbb{S}\right\}.$ (B.8)
By Theorem B.1, $\mathbb{P}_{\mathbf{X}}^{n}(A_{n})\rightarrow 1$ as $n\rightarrow\infty.$
B.4 Proof of Condition (B.3)
@@
for some positive integer $M_{n}\leq T_{\max},$ and will choose $M_{n}$ so that $\mathcal{F}^{n}$ satisfies Condition (B.3) with $\epsilon_{n}^{2}=n^{-{2\alpha\over 2\alpha+d_{\max}}}\log n$. For this purpose, we need the following lemma.
Lemma B.2.
@@
$N(\epsilon_{1},[-n,n]^{T},\|\cdot\|_{1})\leq\bigg(1+{2nT\over\epsilon_{1}}\bigg)^{T}$ (B.9)
@@
$\leq\sum_{t=1}^{T}(1+n)^{|S_{t}|}|\beta_{t}-\tilde{\beta}_{t}|$
$\leq(1+n)^{p}\epsilon_{1},$
where we use the inequality that
$\begin{split}\prod_{j\in S_{t}}\big(\mathbb{I}(x_{j}\leq s_{j})+\mathfrak{a}_{t,j}\mathbb{I}(x_{j}>s_{j})\big)&\leq\prod_{j\in S_{t}}|1+\mathfrak{a}_{t,j}|\\ &\leq(1+n)^{|S_{t}|}.\end{split}$ (B.10)
By letting $\epsilon_{1}=\epsilon/(1+n)^{p}$ in (B.9), we have
$N(\epsilon,\mathcal{F}^{n}(T,\mathcal{E}),\|\cdot\|_{\infty})\leq\bigg(1+{2n(1+n)^{p}T\over\epsilon}\bigg)^{T},$ (B.11)
@@
$N\bigg({\epsilon_{n}\over 36},\mathcal{F}^{n},\|\cdot\|_{\infty}\bigg)$ $\leq\sum_{T=1}^{M_{n}}\sum_{\mathcal{E}\in\mathcal{E}(T)}N\bigg({\epsilon_{n}\over 36},\mathcal{F}^{n}(T,\mathcal{E}),\|\cdot\|_{\infty}\bigg)$
$\leq\sum_{T=1}^{M_{n}}\bigg(1+{72n(1+n)^{p}T\over\epsilon_{n}}\bigg)^{T}\sum_{\mathcal{E}\in\mathcal{E}(T)}1$
$\lesssim\sum_{T=1}^{M_{n}}\bigg(1+{72n(1+n)^{p}T\over\epsilon_{n}}\bigg)^{T}n^{Tp}$ (B.12)
$\leq M_{n}n^{M_{n}p}\bigg(1+{72n(1+n)^{p}M_{n}\over\epsilon_{n}}\bigg)^{M_{n}},$
where the inequality in (B.12) is due to
@@
$\log N\bigg({\epsilon_{n}\over 36},\mathcal{F}^{n},\|\cdot\|_{\infty}\bigg)\lesssim\log M_{n}+M_{n}p\log n+M_{n}\log\bigg(1+{72n(1+n)^{p}M_{n}\over\epsilon_{n}}\bigg)$
With $M_{n}=\lfloor C_{1}{n\epsilon_{n}^{2}\over\log n}\rfloor$ for some large enough constant $C_{1}>0$ (will be determined later), Condition (B.3) is satisfied.
B.5 Proof of Condition (B.4)
For given $\mathbf{x}^{(n)}\in A_{n},$ let $\hat{f}_{S}$ be an ensemble of $t_{n,S}$ many identifiable binary-product trees satisfying $\|f_{0,S}-\hat{f}_{S}\|_{\infty}\leq C_{\mathbb{B}}\epsilon_{n}/2^{p+1},$ whose existence is guaranteed by Theorem B.1, and let $\hat{f}=\sum_{S\in\mathbb{S}}\hat{f}_{S}.$ Let $\hat{\mathcal{E}}_{S}$ and $\hat{\mathcal{B}}_{S}$ be the ensemble partitions and height vector for $\hat{f}_{S}$ and let $\hat{\mathcal{E}}$ and $\hat{\mathcal{B}}$ be the ensemble partition and height vector for $\hat{f}$. Let $t_{n}=\sum_{S\in\mathbb{S}}t_{n,S}$.
@@
$\leq\sum_{t=1}^{t_{n}}(1+n)^{p}|\beta_{t}-\hat{\beta}_{t}|$
$\leq(1+n)^{p}\sqrt{t_{n}}\|\mathcal{B}-\hat{\mathcal{B}}\|_{2},$
where the second inequality follows from (B.10). Hence, if $\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq((1+n)^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2,$ we have
@@
Thus, to prove Condition (B.4), it suffices to show that
$\pi\left\{T=t_{n},\mathcal{E}=\hat{\mathcal{E}},\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq((1+n)^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2\right\}$ (B.13)
is sufficiently large. We decompose (B.13) as
$\pi\left\{T=t_{n},\mathcal{E}=\hat{\mathcal{E}},\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq((1+n)^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2\right\}$
$=\pi\{T=t_{n}\}\pi\{\mathcal{E}=\hat{\mathcal{E}}|T=t_{n}\}\pi\{\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq((1+n)^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2|\mathcal{E}=\hat{\mathcal{E}},T=t_{n}\}.$
@@
B-(a). A lower bound of $\pi\{T=t_{n}\}:$ Since $t_{n}\lesssim n\epsilon_{n}^{2}/\log n$, there exists a constant $d_{2}>0$ such that
$\pi\{T=t_{n}\}$ $={\exp(-C_{*}t_{n}\log n)\over\sum_{t=0}^{T_{\max}}\exp(-C_{*}t\log n)}$
@@
B-(b). A lower bound of $\pi\{\mathcal{E}=\hat{\mathcal{E}}|T=t_{n}\}:$ Let $\mathcal{E}=(\mathcal{R}_{t},t=1,\ldots,t_{n})$ and $\hat{\mathcal{E}}=(\hat{\mathcal{R}}_{t},t=1,\ldots,t_{n}).$ Note that conditional on $T=t_{n},$ $\mathcal{R}_{t},t\in[t_{n}]$ are independent a priori and thus we have
@@
for a certain constant $d_{3}>0.$ B-(c). A lower bound of $\pi\{\|\mathcal{B}-\hat{\mathcal{B}}\|_{2}\leq((1+n)^{p}\sqrt{t_{n}})^{-1}C_{\mathbb{B}}\epsilon_{n}/2|\mathcal{E}=\hat{\mathcal{E}},T=t_{n}\}:$
$\pi\left\{\mathcal{B}\in\mathbb{R}^{t_{n}}:\|\hat{\mathcal{B}}-\mathcal{B}\|_{2}\leq{C_{\mathbb{B}}\over(1+n)^{p}\sqrt{t_{n}}}\times{\epsilon_{n}\over 2}\right\}$
$\geq 2^{-t_{n}}\left({t_{n}\over 2}\right)^{-{t_{n}\over 2}-1}\exp\left(-{\|\hat{\mathcal{B}}\|_{2}^{2}\over\sigma_{\beta}^{2}}-{(C_{\mathbb{B}}\epsilon_{n})^{2}\over 8(1+n)^{2p}t_{n}\sigma^{2}_{\beta}}\right)\left({(C_{\mathbb{B}}\epsilon_{n})^{2}\over 4(1+n)^{2p}t_{n}\sigma_{\beta}^{2}}\right)^{t_{n}\over 2}$ (B.14)
$\geq 2^{-t_{n}}\left({t_{n}\over 2}\right)^{-{t_{n}\over 2}-1}\exp\left(-{t_{n}C^{2}\over\sigma_{\beta}^{2}}-{(C_{\mathbb{B}}\epsilon_{n})^{2}\over 8(1+n)^{2p}t_{n}\sigma_{\beta}^{2}}\right)\left({(C_{\mathbb{B}}\epsilon_{n})^{2}\over 4(1+n)^{2p}t_{n}\sigma_{\beta}^{2}}\right)^{t_{n}\over 2},$ (B.15)
where (B.14) is derived from equation (8.9) in (Ročková and van der Pas (2020)) and (B.15) is derived from Theorem B.1. Finally, the three terms in (B.15) are bounded below by
@@
$\exp\left(-{t_{n}C^{2}\over\sigma_{\beta}^{2}}-{(C_{\mathbb{B}}\epsilon_{n})^{2}\over 8(1+n)^{2p}t_{n}\sigma_{\beta}^{2}}\right)$ $\geq\exp(-d_{5}n\epsilon_{n}^{2})$ (B.16)